# Distribution Aligned Semantics Adaption for Lifelong Person Re-Identification

### v1

@@ Abstract
In real-world scenarios, person Re-IDentification (Re-ID) systems need to be adaptable to changes in space and time. Therefore, the adaptation of Re-ID models to new domains while preserving previously acquired knowledge is crucial, known as Lifelong person Re-IDentification (LReID). Advanced LReID methods rely on replaying exemplars from old domains and applying knowledge distillation in logits with old models. However, due to privacy concerns, retaining previous data is inappropriate. Additionally, the fine-grained and open-set characteristics of Re-ID limit the effectiveness of the distillation paradigm for accumulating knowledge. We argue that a Re-ID model trained on diverse and challenging pedestrian images at a large scale can acquire robust and general human semantic knowledge. These semantics can be readily utilized as shared knowledge for lifelong applications. In this paper, we identify the challenges and discrepancies associated with adapting a pre-trained model to each application domain, and introduce the Distribution Aligned Semantics Adaption (DASA) framework. It efficiently adjusts Batch Normalization (BN) to mitigate interference from data distribution discrepancy and freezes the pre-trained convolutional layers to preserve shared knowledge. Additionally, we propose the lightweight Semantics Adaption (SA) module, which effectively adapts learned semantics to enhance pedestrian representations. Extensive experiments demonstrate the remarkable superiority of our proposed framework over advanced LReID methods, and it exhibits significantly reduced storage consumption. DASA presents a novel and cost-effective perspective on effectively adapting pre-trained models for LReID.
@@ 1 Introduction
Person Re-IDentification (Re-ID) aims at recognizing the same pedestrian across disjoint cameras. With significant advancements in deep learning algorithms over the past decade, it has demonstrated remarkable performance [1, 2, 3, 4]. However, advanced methods usually assume the pedestrian data is provided at once. In real-world scenarios, surveillance data accumulates continuously due to successive deployments and ongoing applications. Re-ID data expands discretely, both spatially and temporally, necessitating the adaptation of Re-ID models to new environments. Consequently, Lifelong person Re-IDentification (LReID) seeks to continuously adapt Re-ID models to novel domains while preserving previously acquired knowledge during the incremental learning process. Unfortunately, due to substantial variations between different Re-ID domains, merely updating the model with new surveillance data can lead to a loss of discriminative ability in previously learned domains. This phenomenon is commonly referred to as the catastrophic forgetting problem.
@@
Recently, we have witnessed the remarkable success of pre-training in computer vision [5, 6, 7], which also revolutionizes the field of person Re-ID. For example, Fu et al. [8, 9] adopt unsupervised pre-training to prevailing Convolutional Neural Network (CNN) architectures like ResNet [10] and achieve significant performance improvement. It is promising to leverage pre-trained models that can be obtained effortlessly for real-world applications. Nevertheless, in the lifelong evolution scenario, directly using the pre-trained weights for model initialization yields limited benefits, as illustrated in Fig. 1 (a). This limitation stems from (1) the inherent drawbacks of the prevailing knowledge distillation paradigm for LReID, and (2) the two crucial techniques in CNN, i.e., Batch Normalization (BN) and Convolutional (Conv) blocks.
@@
More specifically, on the one hand, as depicted in Fig. 1 (b), existing LReID methods [11, 12, 13, 14] follow the distillation baseline [15] to distill knowledge at the logit level. This paradigm heavily relies on the learned classifier. However, unlike conventional lifelong learning [16, 15, 17, 18], LReID is inherently a fine-grained open-set problem. The distillation paradigm would restrict the model’s ability to incorporate valuable knowledge for recognizing unseen identities during inference. It also limits person Re-ID in real-world applications in two aspects. Firstly, although advanced methods that rely on replaying data of old tasks, namely exemplars [19, 20, 21, 22], achieve satisfactory performance, they are impractical for Re-ID applications due to strict privacy constraints associated with pedestrian data. Secondly, the number of classes in the LReID datasets significantly exceeds those in conventional lifelong learning tasks, such as CIFAR-100 [23] and ImageNet [24]. The high dimensional classifier would consume great resources when saving it for knowledge distillation.
On the other hand, blindly tuning the pre-trained Re-ID model [8, 9] in the lifelong learning process would face the problem of catastrophic forgetting. In other words, when adapting the pre-trained model to each application domain, the parameters of BN and Conv blocks should be carefully adjusted. Firstly, BN effectively regularizes the model and captures data distributions, contributing to improved learning [25, 26]. However, it requires careful handling of domain distribution differences between pre-training and application. Secondly, Conv layers specialize in encoding data patterns and semantics at different depths of the network. It is expected that the Conv layers trained on a substantial amount of pedestrian data can acquire robust and generalizable human semantic knowledge [8, 9], thereby serving as a valuable source of shared knowledge across domains. The corresponding problem brought about is how to adapt the acquired general knowledge to each application domain effectively.
To address these challenges, we introduce the Distribution Aligned Semantics Adaption (DASA) framework, which presents a novel LReID paradigm based on pre-trained models. In this paradigm, we do not rely on knowledge distillation that uses saved old data and models for knowledge retention. Instead, we propose freezing the pre-trained Conv layers to preserve shared person semantic knowledge and tuning BN to mitigate domain interference. Additionally, we introduce the lightweight Semantics Adaption (SA) module to effectively aggregate, refine and generate improved pedestrian representations. Our contributions can be summarized as follows:
@@
(1) We advocate taking advantage of the robust and general human semantic knowledge acquired through large-scale pre-training to effectively adapt models in application domains, presenting a new paradigm for LReID.
@@
(2) We propose the Distribution Aligned Semantics Adaption (DASA) framework for LReID. It eliminates the interference of domain distribution discrepancy between pre-training and application, and incorporates the lightweight semantics adaption module to adapt learned semantics for better pedestrian representations.
@@ 2.1 Lifelong Learning
Lifelong learning, also known as incremental learning or continual learning, seeks to maintain stable performance on old tasks while adapting the models to gain new knowledge. Methods can be traditionally divided into three categories, namely regularization-based, rehearsal-based, and architecture-based methods. Regularization-based methods [16, 15] attempt to limit updating important parameters for preceding tasks to mitigate forgetting. To better keep past acquired knowledge, rehearsal-based methods [21, 19, 22] maintain a memory buffer to store finite exemplars of previous tasks. They are prevalent for their satisfactory performance and simplicity. However, rehearsal-based methods generally deteriorate with a smaller buffer size and do not apply to scenarios like Re-ID where data privacy should be considered carefully. Architecture-based methods design separate components or extra parameters for new tasks, so they are immune to forgetting. They expand the network with task-specific components [27, 28, 29] or attend to task-specific sub-networks [30, 31]. For instance, DER [32] dynamically adds a new feature extractor per task, and HAT [31] learns a hard attention mask concurrently to every task. Although lifelong learning has received extensive research in classification or recognition tasks, the presence of imbalanced samples per identity and subtle inter-class variations in person Re-ID poses additional challenges. The significant variations across domains make the problem of catastrophic forgetting even more thorny.
@@ 2.2 Lifelong Person Re-Identification
Recently, there has been remarkable progress in person re-identification leveraging pre-prepared stationary training data [1, 33, 3]. In response to the demand for long-term scenarios, Pu et al. [11] introduce Lifelong person Re-IDentification (LReID) and propose to maintain a learnable knowledge graph to adaptively update previous knowledge. However, due to significant domain variations, it struggles to retain old knowledge without access to previous data. Recently, state-of-the-art LReID methods have drawn inspiration from rehearsal-based lifelong learning approaches. They adopt knowledge distillation baselines designed for traditional lifelong learning to preserve acquired knowledge. For instance, based on the distillation paradigm, GwFReID [14] formulates a comprehensive learning objective for maintaining coherence during progressive learning. PTKP [12] proposes a pseudo task knowledge preservation framework to alleviate the domain gap in the last BN layer. KRC [13] introduces a dynamic memory model for bi-directional knowledge transfers and a knowledge consolidation scheme. However, the intrinsic dissimilarities between Re-ID with the classification task restrict these methods from effectively incorporating and leveraging useful knowledge for LReID. Moreover, privacy concerns limit their applicability in real-world scenarios, where exemplars cannot be stored for lifelong usage.
@@ 3.1 Preliminary
Knowledge distillation paradigm. The serious challenge of diverse variation across different datasets obliges previous LReID methods [14, 13, 12] to save a small number of old exemplars from previous training steps. These exemplars are used for knowledge distillation [15, 21]. In the knowledge distillation paradigm, $\mathcal{G}^{(t-1)}$ and $g^{(t-1)}$ are used to maintain the acquired knowledge. The distillation is performed between the old $\mathcal{G}^{(t-1)}$, $g^{(t-1)}$ and the evolving $\mathcal{G}^{(t)}$, $g^{(t)}$. We denote the extracted feature of an exemplar image $x^{r}$ using $\mathcal{G}^{(t)}$ as ${f^{r}}^{(t)}\in\mathbb{R}^{d}$, and the weight parameter of the classifier $g^{(t)}$ as ${\phi}^{(t)}\in\mathbb{R}^{N^{(t)}\times d}$, where $N^{(t)}=\sum_{j=1}^{t}n^{(j)}_{id}$. The knowledge distillation loss is formulated as follows:
$\mathcal{L}_{kd}=-\frac{1}{|\mathcal{B}|}\sum\limits_{x^{r}\in\mathcal{B}}\sigma\left({f^{r}}^{(t-1)}\cdot{{\phi}^{(t-1)}}^{\top}\right)\ \log\left(\sigma\left({f^{r}}^{(t)}\cdot{{\phi}^{(t)}}^{\top}\right)\right)$ (1)
where $|\mathcal{B}|$ denotes the training batch size, $\sigma$ is the $softmax$ function. For dimension consistency of ${\phi}^{(t-1)}$ and ${\phi}^{(t)}$ in Eq. 1, either only old classes of $g^{(t)}$ are used, or $g^{(t-1)}$ is expanded by using $\mathcal{G}^{(t)}$ to calculate the center of the new classes before the new training step. New samples can also be used for knowledge distillation [15, 11]. Although previous LReID works [12, 13, 14] find it useful to follow the knowledge distillation paradigm, the significant task discrepancy between classification and Re-ID makes them suboptimal. Due to privacy issues, the training data from previous steps should not be available anymore.
Figure 2: The framework of our proposed DASA. The acquired robust and general human semantic knowledge from pre-training is used as shared knowledge, which is kept in the frozen Conv layers. At each training step, we adapt the acquired knowledge from the pre-training domain to the application domain by tuning BN layers and adopting the lightweight Semantics Adaption (SA) modules. During the evolution process of lifelong learning, the previously learned BN and SA are used for initialization in the upcoming domain for forward knowledge transfer, while the old classifier can be freely discarded without increasing the storage burden.
@@ 3.2 Distribution Aligned Semantics Adaption
Distribution alignmnet. Batch Normalization (BN) as a widely-used technique in CNN can normalize the features of each domain to the same reference distribution by domain-specific normalization statistics. It has shown its advantages in regularizing models and improving their generalization ability [25, 26, 34]. Inspired by previous works, we maintain multiple sets of domain-specific BN layers to estimate the feature distribution statistics. We tune BN in each application domain so that the outputs of the updated BN layers exhibit a similar distribution to the pre-training domain, and the remaining frozen pre-trained Conv layers can receive stable input activations. In this way, the learned human semantics can be better used, which improves the discriminability of the model in each application domain.
@@
${\rm BN}(f^{(t)}_{l})={\gamma}_{l}\ \frac{f^{(t)}_{l}-{\mu}^{(t)}_{l}}{\sqrt{{{\sigma}^{(t)}_{l}}^{2}+\epsilon}}+{\beta}_{l}$ (2)
@@
Semantics adaption. Having compensated for the distribution differences between the pre-training domain and the application domain, it is necessary to consider how to efficiently utilize the acquired human semantic knowledge in the application domain. The learned human semantics from pre-training are regarded as robust and general knowledge for Re-ID, so we freeze the pre-trained Conv layers to keep the shared semantic knowledge. Although directly utilizing these general semantics for Re-ID has achieved decent results, it may contribute to inferior performance when the application domain shows great complexity in scenes and pedestrians (as shown in Tab. 6). Therefore, it is necessary to adapt learned general semantics to specific application domains individually.
To this end, we introduce the lightweight Semantics Adaption (SA) module after each frozen Conv layer to aggregate and refine learned semantics. For the sake of efficiency, we implement SA as one depth-wise Conv layer. Assuming $M$ is the channel dimension of the input feature map, the depth-wise Conv layer adopts $M$ kernels and each for one channel of the input feature map, respectively. One of the direct designs is using the 1$\times$1 depth-wise Conv layer. However, we find it has no effect on performance improvement and even deteriorates the discriminative ability of the model (as shown in Fig. 5). It makes sense since the spatial information is not taken into consideration. The human body structure and the relations between different body parts can be regarded as one of general human semantics. Therefore, we propose to apply a relatively large kernel size to aggregate the general semantics to represent pedestrians effectively. Note that since the SA module performs convolutional operation at each channel separately, the increase in kernel size would not result in a significant increase in parameters. In our experiments, we find that the kernel size of 5$\times$5 is able to achieve a good balance between effectiveness and efficiency.
After training in each domain, we can obtain a set of domain-specific SA modules, i.e., $SA^{(t)}={\left\{w^{(t)}_{l}\right\}}^{L_{Conv}}_{l=1}$, where $L_{Conv}$ is the number of Conv layers in the CNN model. Similarly to our distribution alignment design, to achieve the forward transfer of knowledge from the learned domain to the next, we use the $SA^{(t-1)}$ as the initial to train $SA^{(t)}$ effectively.

### v2

@@ Abstract
In real-world scenarios, person Re-IDentification (Re-ID) systems need to be adaptable to changes in space and time. Therefore, the adaptation of Re-ID models to new domains while preserving previously acquired knowledge is crucial, known as Lifelong person Re-IDentification (LReID). Advanced LReID methods rely on replaying exemplars from old domains and applying knowledge distillation in logits with old models. However, due to privacy concerns, retaining previous data is inappropriate. Additionally, the fine-grained and open-set characteristics of Re-ID limit the effectiveness of the distillation paradigm for accumulating knowledge. We argue that a Re-ID model trained on diverse and challenging pedestrian images at a large scale can acquire robust and general human semantic knowledge. These semantics can be readily utilized as shared knowledge for lifelong applications. In this paper, we identify the challenges and discrepancies associated with adapting a pre-trained model to each application domain and introduce the Distribution Aligned Semantics Adaption (DASA) framework. It efficiently adjusts Batch Normalization (BN) to mitigate interference from data distribution discrepancy and freezes the pre-trained convolutional layers to preserve shared knowledge. Additionally, we propose the lightweight Semantics Adaption (SA) module, which effectively adapts learned semantics to enhance pedestrian representations. Extensive experiments demonstrate the remarkable superiority of our proposed framework over advanced LReID methods, and it exhibits significantly reduced storage consumption. DASA presents a novel and cost-effective perspective on effectively adapting pre-trained models for LReID. The code is available at https://github.com/QizaoWang/DASA-LReID.
@@ 1 Introduction
Person Re-IDentification (Re-ID) aims at recognizing the same pedestrian across disjoint cameras. With significant advancements in deep learning algorithms over the past decade, it has demonstrated remarkable performance [1, 2, 3, 4, 5, 6, 7]. However, advanced methods usually assume the pedestrian data is provided at once. In real-world scenarios, surveillance data accumulates continuously due to successive deployments and ongoing applications. Re-ID data expands discretely, both spatially and temporally, necessitating the adaptation of Re-ID models to new environments. Consequently, Lifelong person Re-IDentification (LReID) seeks to continuously adapt Re-ID models to novel domains while preserving previously acquired knowledge during the incremental learning process. Unfortunately, due to substantial variations between different Re-ID domains, merely updating the model with new surveillance data can lead to a loss of discriminative ability in previously learned domains. This phenomenon is commonly referred to as the catastrophic forgetting problem.
@@
Recently, we have witnessed the remarkable success of pre-training in computer vision [8, 9], which also revolutionizes the field of person Re-ID. For example, Fu et al. [10, 11] adopt unsupervised pre-training to prevailing Convolutional Neural Network (CNN) architectures like ResNet [12] and achieve significant performance improvement. It is promising to leverage pre-trained models that can be obtained effortlessly for real-world applications. Nevertheless, as observed in Fig. 1 (a), directly using the pre-trained weights for model initialization yields limited benefits in the lifelong evolution scenario. Advanced LReID methods [13, 14, 15, 16] follow the prevailing knowledge distillation paradigm [17]. Thus we are motivated to think whether the performance degradation stems from the inherent drawbacks of the knowledge distillation paradigm.
@@
As depicted in Fig. 1 (b), existing LReID methods follow the distillation baseline to distill knowledge at the logit level. This paradigm heavily relies on the learned classifier. However, unlike conventional lifelong learning [18, 17, 19, 20], LReID is inherently a fine-grained open-set problem. The distillation paradigm would restrict the model’s ability to incorporate valuable knowledge for recognizing unseen identities during inference. It also limits person Re-ID in real-world applications in two aspects. Firstly, relying on replaying data of old tasks, namely exemplars [21, 22, 23, 24], is impractical due to strict privacy constraints associated with pedestrian data. Secondly, the number of classes in the LReID datasets significantly exceeds those in conventional lifelong learning tasks, such as ImageNet [25]. The high dimensional classifier would consume great resources when saving it for knowledge distillation.
Therefore, to achieve a balance between knowledge preservation and updating with minimal source consumption, and without using exemplars, we introduce the Distribution Aligned Semantics Adaption (DASA) framework, which presents a novel LReID paradigm based on pre-trained models. In this paradigm, we do not rely on knowledge distillation that uses saved old data and models for knowledge retention. Instead, we exert the advantages of the pre-trained model in the lifelong learning process. Specifically, we carefully handle and leverage the two crucial techniques in CNN, i.e., Batch Normalization (BN) and Convolutional (Conv) blocks. Firstly, BN effectively regularizes the model and captures data distributions, contributing to improved learning [26, 27]. In our proposed paradigm, we tune BN effectively to deal with domain distribution differences between pre-training and application. Secondly, Conv layers specialize in encoding data patterns and semantics at different depths of the network. It is expected that the Conv layers trained on a substantial amount of pedestrian data can acquire robust and generalizable human semantic knowledge [10, 11], thereby serving as a valuable source of shared knowledge across domains. Therefore, we propose freezing the pre-trained Conv layers to preserve shared person semantic knowledge. However, general semantics are not capable of distinguishing pedestrians in complex application scenarios. To adapt the acquired general knowledge to each application domain effectively, we introduce the lightweight Semantics Adaption (SA) module, which can efficiently aggregate, refine, and generate improved pedestrian representations. Adapting the general human semantics to more representative identity semantics represents a process from coarse to fine, showing great ability to evolve the model for lifelong application. Our contributions can be summarized as follows:
@@
(1) We advocate taking advantage of the robust and general human semantic knowledge acquired through large-scale pre-training to effectively adapt models in application domains, presenting a new paradigm free of exemplars for LReID.
@@
(2) We propose the Distribution Aligned Semantics Adaption (DASA) framework for LReID. It eliminates the interference of domain distribution discrepancy between pre-training and application, and incorporates the lightweight semantics adaption module to aggregate and refine learned semantics for better pedestrian representations.
@@ 2.1 Lifelong Learning
Lifelong learning, also known as incremental or continual learning, seeks to maintain stable performance on old tasks while adapting the model to gain new knowledge. Methods can be traditionally divided into three categories, namely regularization-based, rehearsal-based, and architecture-based methods. Regularization-based methods [18, 17] limit updating important parameters for preceding tasks to mitigate forgetting. To better keep past acquired knowledge, rehearsal-based methods [23, 21, 24] maintain a memory buffer to store finite exemplars of previous tasks. They are prevalent for their satisfactory performance and simplicity. However, rehearsal-based methods generally deteriorate with a smaller buffer size and do not apply to scenarios where data privacy should be considered carefully. Architecture-based methods design separate components or extra parameters for new tasks, so they are immune to forgetting. They expand the network with task-specific components [28, 29] or attend to task-specific sub-networks [30, 31]. For instance, DER [32] dynamically adds a new feature extractor per task, and HAT [31] learns a hard attention mask concurrently to every task. Although lifelong learning has received extensive research in classification or recognition tasks, the presence of imbalanced samples per identity and subtle inter-class variations in person Re-ID poses additional challenges. The significant variations across domains make the problem of catastrophic forgetting even more thorny.
@@ 2.2 Lifelong Person Re-Identification
Recently, there has been remarkable progress in person re-identification leveraging pre-prepared stationary training data [1, 33, 3, 5, 34, 35]. In response to the demand for long-term scenarios, Pu et al. [13] introduce Lifelong person Re-IDentification (LReID) and propose to maintain a learnable knowledge graph to adaptively update previous knowledge. However, due to significant domain variations, it struggles to retain old knowledge without access to previous data. Recently, state-of-the-art LReID methods have drawn inspiration from rehearsal-based lifelong learning approaches. They adopt knowledge distillation baselines designed for traditional lifelong learning to preserve acquired knowledge. For instance, based on the distillation paradigm, GwFReID [16] formulates a comprehensive learning objective for maintaining coherence during progressive learning. PTKP [14] proposes a pseudo task knowledge preservation framework to alleviate the domain gap in the last BN layer. KRC [15] introduces a dynamic memory model for bi-directional knowledge transfers and a knowledge consolidation scheme. However, the intrinsic dissimilarities between Re-ID with the classification task restrict these methods from effectively incorporating and leveraging useful knowledge for LReID. Moreover, privacy concerns limit their applicability in real-world scenarios, where exemplars cannot be stored for lifelong usage. The recent work Teata [36] achieves knowledge alignment, transfer, and accumulation within an “image-text-image” closed loop, leveraging the advantages of text semantics while avoiding reliance on exemplars.
In LReID, we aim to achieve a balance between knowledge preservation and updating with minimal source consumption, and without using exemplars. It differs from existing studies in Test-Time Adaption (TTA) [37, 38] and Domain Generation (DG) [39, 40]. Specifically, in TTA, the pre-trained model is optimized during inference [37]. For instance, BNTA [38] adapts the model using gallery data and self-supervised auxiliary tasks, and TEMP [41] uses query images and gallery features to minimize entropy at test time. Differently, in LReID, the model is updated during training and directly tested on seen and unseen domains. In DG, studies aim to improve the generalization performance of the model on unseen domains [39, 40], and the model is usually trained on multiple domains to learn domain-invariant representations. In contrast, in LReID, the model is optimized with various domains sequentially and the data from previous domains is inaccessible, so knowledge forgetting will be a great challenge.
@@ 3.1 Preliminary
Knowledge distillation paradigm. The serious challenge of diverse variation across different datasets obliges previous LReID methods [16, 15, 14] to save a small number of exemplars from previous training steps and use them for knowledge distillation [17, 23]. In the knowledge distillation paradigm, $\mathcal{G}^{(t-1)}$ and $g^{(t-1)}$ are used to maintain the acquired knowledge. The distillation is performed between the old $\mathcal{G}^{(t-1)}$, $g^{(t-1)}$ and the evolving $\mathcal{G}^{(t)}$, $g^{(t)}$. We denote the extracted feature of an exemplar image $x^{r}$ using $\mathcal{G}^{(t)}$ as ${f^{r}}^{(t)}\in\mathbb{R}^{d}$, and the weight parameter of the classifier $g^{(t)}$ as ${\phi}^{(t)}\in\mathbb{R}^{N^{(t)}\times d}$, where $N^{(t)}=\sum_{j=1}^{t}n^{(j)}_{id}$. The knowledge distillation loss is formulated as follows:
$\mathcal{L}_{kd}=-\frac{1}{|\mathcal{B}|}\sum\limits_{x^{r}\in\mathcal{B}}\sigma\big{(}{f^{r}}^{(t-1)}\cdot{{\phi}^{(t-1)}}^{\top}\big{)}\ \log\left(\sigma\big{(}{f^{r}}^{(t)}\cdot{{\phi}^{(t)}}^{\top}\big{)}\right),$ (1)
where $|\mathcal{B}|$ denotes the training batch size, $\sigma$ is the $softmax$ function. For dimension consistency of ${\phi}^{(t-1)}$ and ${\phi}^{(t)}$ in Eq. 1, either only old classes of $g^{(t)}$ are used, or $g^{(t-1)}$ is expanded by using $\mathcal{G}^{(t)}$ to calculate the center of the new classes before the new training step. New samples can also be used for distillation [17, 13]. Previous LReID works [14, 15, 16] all follow the knowledge distillation paradigm, but the significant task discrepancy between classification and Re-ID makes them suboptimal. Due to privacy issues, the training data from previous steps should not be available anymore.
Figure 2: The framework of DASA. The acquired robust and general human semantics from pre-training are used as shared knowledge, which is kept in the frozen Conv layers. At each training step, we adapt the acquired knowledge from the pre-training to application domain by tuning BN layers and adopting the lightweight Semantics Adaption (SA) modules. During the lifelong evolution process, the previously learned BN and SA are used for initialization in the upcoming domain for forward knowledge transfer, while the old classifier can be discarded without increasing storage burden.
@@ 3.2 Distribution Aligned Semantics Adaption
Distribution alignment. Batch Normalization (BN) as a widely-used technique in CNN can normalize the features of each domain to the same reference distribution by domain-specific normalization statistics. It has shown its advantages in regularizing models and improving their generalization ability [27, 42]. Inspired by previous works, we maintain multiple sets of domain-specific BN layers to estimate the feature distribution statistics. We tune BN in each application domain so that the outputs of the updated BN layers exhibit a similar distribution to the pre-training domain, and the remaining frozen pre-trained Conv layers can receive stable input activations. In this way, the learned human semantics can be better used, which improves the discriminability of the model in each application domain.
@@
${\rm BN}(f^{(t)}_{l})={\gamma}_{l}\ \frac{f^{(t)}_{l}-{\mu}^{(t)}_{l}}{\sqrt{{{\sigma}^{(t)}_{l}}^{2}+\epsilon}}+{\beta}_{l},$ (2)
@@
Semantics adaption. Having compensated for the distribution differences between the pre-training domain and the application domain, it is time to consider how to efficiently utilize the acquired human semantic knowledge in the application domain. The learned human semantics from pre-training are regarded as robust and general knowledge for Re-ID, so we freeze the pre-trained Conv layers to keep the shared knowledge. Although directly utilizing these general semantics for Re-ID has achieved decent results, it may contribute to inferior performance when the application domain shows great complexity in scenes and pedestrians (as shown in Tab. 6). Therefore, it is necessary to adapt the learned general semantics to specific application domains.
To this end, we introduce the lightweight Semantics Adaption (SA) module after each frozen Conv layer to aggregate and refine learned semantics effectively. For the sake of efficiency, it could be a good choice to implement SA as one depth-wise Conv layer. Assuming $M$ is the channel dimension of the input feature map, the depth-wise Conv layer adopts $M$ kernels and each for one channel of the input feature map, respectively. One of the direct designs is using the 1 $\times$ 1 depth-wise Conv layer. However, we find it has no effect on performance improvement and even deteriorates the discriminative ability of the model (as shown in Fig. 5). It makes sense since there is no interaction in the spatial dimension to achieve semantics adaption. Intuitively, the information interaction across spatial is vital to capture discriminative semantics, such as body figures, and improve the robustness of learned semantics in complex scenes. Therefore, we propose to apply a relatively large kernel size to aggregate and refine the general semantics. Note that since the SA module performs convolutional operation at each channel separately, the increase in kernel size would not result in a significant increase in parameters. In our experiments, we find that the kernel size of 5 $\times$ 5 can achieve a good balance between effectiveness and efficiency.
There are some differences when comparing the SA module with LoRA [43], which adds trainable pairs of rank decomposition matrices in parallel to existing weight matrices. Technically, LoRA updates the original weights and changes the calculation operator, adapting the model for different tasks. Differently, the SA module aims to leverage general human semantics, improving the Re-ID discriminative ability of the model in different application domains. Adapting the general human semantics to more representative identity semantics represents a process from coarse to fine, showing great ability to evolve the model for lifelong application. Incorporating reparametrization techniques like LoRA [43] can be a prospective future work.
After training in each domain, we can obtain a set of domain-specific SA modules, i.e., $SA^{(t)}={\left\{w^{(t)}_{l}\right\}}^{L_{Conv}}_{l=1}$, where $w^{(t)}_{l}$ denotes the parameters of the SA module at layer $l$, and $L_{Conv}$ is the number of Conv layers in the CNN model. Similarly to our distribution alignment design, to achieve the forward transfer of knowledge from the learned domain to the next, we use the $SA^{(t-1)}$ as the initial to train $SA^{(t)}$ effectively.